The perimeter of a triangle is 60 inches. The second side is 10 inches longer than the first side. The third side is 6 inches more than twice the first side. what is the length of the smallest side?
Answer by usyim88hk(157)

Let x be the length of the first side:
x = first side
x+10 = second side
2x+6 = third side
Now set up an equation:
x+x+10+2x+6 = 60 (because three sides add up to 60)
4x + 16 = 60
4x = 44
x = 11
So the length of the first side is 11
the length of the second side is 21
the length of the third side is 28